solve:
3x-4y=26
{
x-8y=22

Answer by ewatrrr(24785)

 
Hi,
solve:
3x-4y=26
x-8y=22
3x-4y =26
-3x+24y=-66 |multiplying thru by -3 t eliminate the x variable
20y = -40
y = -2 and x = 6 ( x = 8*-2 + 22)
CHECKING our Answer***
3x-4y=26
18 + 8 = 26